OpenStudy (geometry_hater):

I don't know what this is talking about Find the vaule of a b and c that will garuntee that the system has exactly one solution x+2y = 3 ax+by = c

OpenStudy (geometry_hater):

howd you do that?

myininaya (myininaya):

first I'm going to put both in y=mx+b form \[y=\frac{-x}{2}+\frac{3}{2}\] \[y=\frac{-ax}{b}+\frac{c}{b}\] Parallel lines would give us no solution. Parallel lines are lines with same slope and different y-intercepts. Same lines would give us infinitely many lines. These are lines with same slope and same y-intercept. Any other types would give us one intersection. So this is what we want: \[\frac{-1}{2} \neq \frac{-a}{b} \]

myininaya (myininaya):

For example if a =5 and b=2 and c= whatever like 1 We have what we want above \[\frac{-1}{2} \neq \frac{-5}{2}\] so we would have the following two lines \[y=\frac{-x}{2}+\frac{3}{2}\] \[y=\frac{-5x}{2}+\frac{1}{2}\] To show you that this does have one intersection I will solve the system \[\frac{-x}{2}+\frac{3}{2}=\frac{-5x}{2}+\frac{1}{2}\] first I multiply both sides by 2 \[-x+3=-5x+1\] \[5x-x=1-3 =>4x=-2 => x=\frac{-2}{4}=\frac{-1}{2}=> y=\frac{\frac{-1}{2}}{2}+\frac{3}{2}=\frac{-1}{4}+\frac{6}{4}=\frac{5}{4}\] This is just one possible set of values that a,b,c can be

OpenStudy (geometry_hater):

wow ok thanks

myininaya (myininaya):

you just don't want -1/2 to equal -a/b

myininaya (myininaya):

so choosing a=2 and b=4 would be wrong do you see why?

OpenStudy (geometry_hater):

yes

myininaya (myininaya):

because 2/4 reduces to 1/2

myininaya (myininaya):

they are the same fraction